A company paid dividend $12. Company is expected to grow the dividend at 10% per year for next 3 years and then at 6% per year forever. What is the price of stock at the end of third year if your discount rate is 14%? 


1
Expert's answer
2021-05-17T08:06:50-0400

After-tax yield = Rate on the taxable bond x ( Tax rate-1) ... Your annual dividend income on $12 shares would be total of rate amount(14+10+6)% x $12======> 30/100×12 =(0.3×12)

= 3.6×( tax rate -1)

========> 3.6×(3-1) = (3.6×2)= $7.2

Therefore, it will be $7.2 price of stock at the end of third year if the discount rate is 14%